These are the hourly salaries od all workers employed at the Manhattan Law Firm together with the probabilities

hourly salary in dollars 8 90 125 180 350
probability (distribution) of having such a salary .1 .2 .5 .1 .1

The standard deviation for this salary distribution is approximately
Answer #1

Answer

we have the following data

x P(x)
8 0.1
90 0.2
125 0.5
180 0.1
350 0.1

First, we need to find the mean

we know mean = r P(z)

setting the values, we get

Mean= μ (8 * 0.1) + (90 * 0.2) + (125 * 0.5) + (180 * 0.1) + (350 * 0.1) = 134.3

Now, the formula for standard deviation is given as

sigma= sqrt{[sum (x^2*P(x))] - mu^2}

setting the given values, we get

νΊδ2 * 0.1) + (902 * 0.2) + (125? * 0.5) + (1802 * 0.1) (3502 * 0.1)]-(134.3)2 σ

on solving, we get

σ = V/24928,9| _ 18036.49 = V689241-83.02

So, the required standard deviation for this salary distribution is approximately is 83.02 (rounded to 2 decimals)